OpenStudy (anonymous):

Let g(x) be the reflection of f(x) = 3x + 2 in the x-axis. What is a function rule for g(x)? g(x) = 3x – 2 g(x) = –3x – 2 g(x) = 3x + 2 g(x) = –3x + 2

OpenStudy (dumbcow):

a reflection of f(x) on x-axis = -f(x) = -(3x+2) = -3x -2

OpenStudy (anonymous):

thank you . .

OpenStudy (dumbcow):

whereas reflection of f(x) on y-axis would be f(-x) = 3(-x) +2 = -3x+2

OpenStudy (dumbcow):

yw